. Given the following system of equations
x + y = 4k
2x = k
where k is a constant real number, what method would you use to solve this system and
why? Solve the system for (x, y)in terms of k . List your solution as an ordered pair.

Answer by stanbon(75887)

Given the following system of equations
x + y = 4k
2x = k
where k is a constant real number, what method would you use to solve this system and
why? Solve the system for (x, y)in terms of k . List your solution as an ordered pair.
-----------------
x = k/2
Substitute into the 1st equation to solve for "y":
(k/2) + y = 4k
y = (7/2)k
---------------------
Solution: (k/2 , 7k/2)
============================
Cheers,
Stan H.